# Core Safety Values for Provably Corrigible Agents

### v1

@@ Appendix G Proof of Proposition 5
Our object of study will be the sensitivity‑1 Boolean query $\phi_{H}=\mathbf{1}\!\left[B\text{ occurs by time }H\right]$. For this indicator, the Laplace mechanism in the centralized setting (Dwork and Roth 2014, Def. 3.3) outputs $\widetilde{\phi}_{H}=\phi_{H}+\eta$, where $\eta$ is Laplace noise with $\mathbb{E}[\eta]=0$ and $\mathrm{Var}[\eta]=2\varepsilon^{-2}$ (scale $1/\varepsilon$). The Laplace mechanism is $\left\langle\varepsilon,0\right\rangle$-DP (Dwork and Roth 2014, Theorem 3.6); therefore it also satisfies $\left\langle\varepsilon,\delta_{\mathrm{priv}}\right\rangle$-DP for any $0<\delta_{\mathrm{priv}}<1$. Furthermore, each user can run an $\varepsilon$-local Laplace randomizer with the same scale (or, more commonly, randomized response), and this would be $\left\langle\varepsilon,0\right\rangle$-DP as well (Dwork and Roth 2014, Obs. 12.1).
Balcan et al. (2012, Theorem 15) show that for a statistical query answered from a sample of size $N$, Laplace noise of scale $b_{\mathrm{BBFM}}=O\left(\frac{\sqrt{\log(1/\delta_{\mathrm{priv}})}}{\varepsilon\sqrt{N}}\right)$ achieves $\left\langle\varepsilon,\delta_{\mathrm{priv}}\right\rangle$-DistP. We will show below that a sample size of
@@
Furthermore, since each sample requires at most $H$ simulator steps, the total time is $\operatorname{poly}(n,\varepsilon^{-1},\log(1/\delta))$ by the assumed bound on $H$; hence the algorithm is in $\mathsf{BPP}$, by construction. Since the verifier alone can decide membership in time $\mathsf{BPP}$, we may invoke the classic Goldreich and Krawczyk (1996) compiler, which converts any $\mathsf{BPP}$ decision procedure into a constant‑round, public‑coin statistical zero‑knowledge ($\mathsf{SZK}$) protocol. Concretely, in our setting the prover sends no messages: the verifier simply publishes its random tape and reruns the $\mathsf{BPP}$ test locally. Because the entire interaction transcript consists solely of this public randomness, a simulator can reproduce the transcript exactly (it re‑samples the same random tape), so the verifier’s view in the real and simulated executions is identical. Hence the empty‑transcript protocol is SZK, and the language $\mathsf{SAFE}^{\mathrm{priv}}_{H,\varepsilon,\delta}$ belongs to $\mathsf{SZK}$ as claimed. ∎

### v2

@@ Appendix H Proof of Proposition 5
Our object of study will be the sensitivity‑1 Boolean query $\phi_{H}=\mathbf{1}\!\left[B\text{ occurs by time }H\right]$. For this indicator, the Laplace mechanism in the centralized setting (dwork2014algorithmic, Def. 3.3) outputs $\widetilde{\phi}_{H}=\phi_{H}+\eta$, where $\eta$ is Laplace noise with $\mathbb{E}[\eta]=0$ and $\mathrm{Var}[\eta]=2\varepsilon^{-2}$ (scale $1/\varepsilon$). The Laplace mechanism is $\left\langle\varepsilon,0\right\rangle$-DP (dwork2014algorithmic, Theorem 3.6); therefore it also satisfies $\left\langle\varepsilon,\delta_{\mathrm{priv}}\right\rangle$-DP for any $0<\delta_{\mathrm{priv}}<1$. Furthermore, each user can run an $\varepsilon$-local Laplace randomizer with the same scale (or, more commonly, randomized response), and this would be $\left\langle\varepsilon,0\right\rangle$-DP as well (dwork2014algorithmic, Obs. 12.1).
balcan2012distributed show that for a statistical query answered from a sample of size $N$, Laplace noise of scale $b_{\mathrm{BBFM}}=O\left(\frac{\sqrt{\log(1/\delta_{\mathrm{priv}})}}{\varepsilon\sqrt{N}}\right)$ achieves $\left\langle\varepsilon,\delta_{\mathrm{priv}}\right\rangle$-DistP. We will show below that a sample size of
@@
Furthermore, since each sample requires at most $H$ simulator steps, the total time is $\operatorname{poly}(n,\varepsilon^{-1},\log(1/\delta))$ by the assumed bound on $H$; hence the algorithm is in $\mathsf{BPP}$, by construction. Since the verifier alone can decide membership in time $\mathsf{BPP}$, we may invoke the classic goldreich1996composition compiler, which converts any $\mathsf{BPP}$ decision procedure into a constant‑round, public‑coin statistical zero‑knowledge ($\mathsf{SZK}$) protocol. Concretely, in our setting the prover sends no messages: the verifier simply publishes its random tape and reruns the $\mathsf{BPP}$ test locally. Because the entire interaction transcript consists solely of this public randomness, a simulator can reproduce the transcript exactly (it re‑samples the same random tape), so the verifier’s view in the real and simulated executions is identical. Hence the empty‑transcript protocol is $\mathsf{SZK}$, and the language $\mathsf{SAFE}^{\mathrm{priv}}_{H,\varepsilon,\delta}$ belongs to $\mathsf{SZK}$ as claimed. ∎